OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4a + 20b = 8 8a = -40b + 16

OpenStudy (anonymous):

Substitution means you substitute one variable in for another. So in this case we can solve for a: first equation: a = (8-20b)/4 now SUBTITUTE that into the second equation: 8[(8-20b)/4] = -40b + 16 and solve for b.

hero (hero):

8a = -40b + 16 8a + 40b = 16 2(4a + 20b) = 16 4a + 20b = 16/2 4a + 20b = 8 Thus 4a + 20b = 8 is equivalent to 8a = -40b + 16, which means they are the same equation.

OpenStudy (anonymous):

Thanks tons! This substitution stuff is kicking my butt! I appreciate all the help!

hero (hero):

yw =]

hero (hero):

Substitution is just playing with math. I can show you a few tricks

OpenStudy (anonymous):

please!!

hero (hero):

We can go on vyew and I can show you.

OpenStudy (anonymous):

what is that?

hero (hero):

It's a site with an online whiteboard and chat amongst other things. Great for collaboration.

OpenStudy (anonymous):

ok. let me go look.

hero (hero):

Let me set it up. Give me a minute.